# Robust Self-calibration of Focal Lengths from the Fundamental Matrix

### v2

@@ 3.2 Real Focal Length Checking within RANSAC
The method proposed in 3.1 is usually applied on a fundamental matrix $\mathbf{F}$ that is obtained using a robust RANSAC-style estimation [11, 1, 8, 9, 16, 7]. Compared to the Bougnoux formula [2], our method does not suffer from imaginary focal length estimates in the presence of noise or an error in the principal point location. Thus, our method does not need to check whether the fundamental matrix returned by the RANSAC is decomposible to real focal lengths or not. Nevertheless, as we observed in our experiments, fundamental matrices that result in imaginary focal lengths usually do not provide good estimates and are therefore not selected inside RANSAC as best models.
@@
To reject models, we use a modified version of the Bougnoux formula [2] presented in [32], which has the form of a ratio of two polynomials in the elements of the fundamental matrix. The formula returns squared focal lengths. When the sign of one of the squared focal lengths is negative, we reject the model. This approach does not require the use of SVD or other expensive matrix manipulations. This type of checking can be implemented into different variants of RANSACs and, as shown later in our experiments, it leads to reduced computational times while maintaining or even improving accuracy of pose estimates. Implementation details can be found in the Supplementary Material (SM).
@@ 4 Synthetic Experiments
(c) Phototourism [17]
Figure 1: Plots showing portion of samples for which our algorithm would converge given a threshold for the relative change of errors in successive iterations $\frac{|e^{n}-e^{n-1}|}{e^{n}}<\epsilon$. For synthetic experiments (a) and (b) we generated 1000 samples with added noise ($\sigma_{n}=1$, $\sigma_{p}=10$). We set the priors as $f_{1}^{p}=660,f_{2}^{p}=440$. For (c) we used the Phototourism benchmark dataset [17]. See Section 4 for details.
@@
In all our synthetic and real experiments, we set the weights in (2) to $w_{i}^{f}=5\cdot 10^{-4}$ and $w_{i}^{\mathbf{c}}=1.0$. These weights were obtained by testing different combinations on the Brandenburg Gate validation scene from the Phototurism dataset [17] and resulted in the best performance on this scene. However, as we show also in synthetic experiments, for a wide variety of combinations of weights, we consistently obtain very good estimates and outperform the state-of-the-art methods. We set the weights for Hartley and Silpa-Anan’s method [14] in the same manner to 1.0 for $w_{F}$, $10^{-4}$ for $w_{\mathbf{c}}$ and $10^{-6}$ for $w_{f}$.
@@
To perform the experiments, we set up a scene with two cameras. We assume $640~{}\times~{}480$ image size and principal points in image centers. The first camera has focal length $f_{1}=600$ and the second $f_{2}=400$. The first camera has center at the origin with the z-axis of the world coordinate system being its principal axis. We perform experiments with the second camera in different configurations denoted as $\mathcal{C}(\theta,y)$. For $\mathcal{C}(0\degree,y)$ the second camera center is located at coordinates $(1200,y,600)$. The camera is rotated by $60\degree$ around its $y$-axis. With $\theta\neq 0\degree$ we additionally rotate the second camera around the $x$-axis by $\theta$. When $y=0$ and $\theta=0\degree$ the principal axes of the two cameras intersect and are thus coplanar, which is a degenerate configuration for the Bougnoux formula [2].
For a given camera setup, we uniformly generate 100 random points visible by both cameras. After projecting the points into both images, we add Gaussian noise with a standard deviation $\sigma_{n}$ to the pixel coordinates. We also emulate errors in the assumed locations of the principal points. The principal points are shifted from the image centers by a distance sampled from the normal distribution with a standard deviation of $\sigma_{p}$. We estimate the fundamental matrix using MAGSAC++ [1] as implemented in OpenCV [3].
@@
In the first experiment, we evaluate convergence of the proposed iterative method. Different termination criteria can be considered. We terminate the algorithm when the relative change in error in two successive iterations $\frac{|e^{n}-e^{n-1}|}{e^{n}}$ is less than a predetermined threshold. Fig. 1(a) shows that for non-degenerate configuration our method converges quickly even with strict thresholds. Fig. 1(b) shows the convergence rates of our method when the principal axes are coplanar. The degenerate configuration requires an increased number of iterations to converge with the same thresholds. However, even in a degenerate configuration, the method converges in a few iterations for most of the samples, even with strict thresholds. We also evaluate convergence on a large scale real-world dataset [17] (see section 4 for more details). As shown in Fig. 1(c), our method successfully converges for most of the samples even with a strict threshold. Using these findings, we established the maximum number of iterations for our method at 50, as further iterations demonstrated diminishing returns in terms of convergence rate.
@@
In the next experiment, we compare our method with state-of-the-art methods in terms of accuracy of the estimated focal lengths. The methods considered are the baseline Bougnoux formula [2] and the iterative approaches of Hartley and Silpa-Anan [14] and Fetzer et al. [10]. In the experiments, we use the same priors for Hartley and Silpa-Anan’s method as for ours. We set the initial values for LM optimization for the method by Fetzer et al. to the same values as our priors.
OursHartley and Silpa-Anan[14]Fetzer et al. [10]Bougnoux[2]GT $f_{1}$rior $f_{1}$
@@ 5 Real-world Experiments
Figure 3: The plot shows mAAp achieved on the Phototourism datset [17] for different RANSAC implementations [3, 16, 24] with and without real focal length checking (RFC) under varying number of iterations. To calculate the poses we use the ground truth focal lengths. Further details are provided in the SM.
OursHartley and Silpa-Anan[14]Fetzer et al. [10]Bougnoux[2]Sturm [41]Minimal 6-point [40]PriorRFCOriginal
(a) Phototourism [17]
(b) Aachen Day-Night v1.1 [47]
(c) ETH3D Multiview [37]
@@
Phototourism [17] Aachen Day-Night v1.1 [47]
@@
Table 1: Median errors for poses ($p_{err}$) and focal lengths ($f_{i}^{err}$) and mean average accuracy scores for poses (mAAp) and estimated focal lengths (mAAf) on 12 scenes from the Phototourism dataset [17] and the Aachen Day-Night v1.1 dataset [47]. RFC denotes real focal length checking as described in subsection 3.2.
@@
We assess the real-world performance of our method using two extensive datasets. The first dataset, referred to as the Phototourism dataset, was originally introduced in [17] to serve as a robust benchmark to evaluate structure-from-motion pipelines. The dataset contains 25 scenes of landmarks. Large sets of crowd-sourced images are available for each scene. For 13444Originally, reconstructions were available for 16 of the scenes, but 3 of the scenes were later dropped from the dataset due to data inconsistencies. of the scenes a COLMAP [36] reconstruction is provided to serve as ground truth for camera intrinsics and extrinsics. To evaluate the estimated focal lengths, we randomly sample 1000 pairs of images for each scene. We only consider pairs with sufficient co-visibility as defined in [17]. We used the Brandenburg Gate scene to serve as a validation set to determine the optimal setting of the weights $w_{i}^{f}$ and $w_{i}^{\mathbf{c}}$ in (2).
The second dataset, known as the Aachen Day-Night v1.1 Dataset, is an extension of the previously established Aachen Day-Night dataset [34, 35]. This dataset incorporates a reference COLMAP [36] model that has been reconstructed from a comprehensive collection of 6,697 images captured by a variety of cameras and smartphones. For evaluation, we randomly sample 1000 pairs of images that adhere to the same co-visibility criteria as those defined for the Phototourism dataset.
@@
We use LoFTR [43] to obtain point correspondences. We perform inference on images resized so that the largest dimension is equal to 1024 px. To estimate the fundamental matrix, we use MAGSAC++ [1] with the epipolar threshold set to 3 px.
We also perform experiments with and without utilizing real focal length checking (RFC) as described in subsection 3.2. Fig 3 shows that on the Phototourism dataset [17] the fundamental matrices obtained with RFC generally lead to more accurate pose estimates while also significantly reducing computational times across various RANSAC implementations. Implementation details and further experiments are provided in SM.
@@
We compare our method with three state-of-the-art approaches. We evaluate the Bougnoux formula [2], Hartley and Silpa-Anan’s iterative method [14] and the iterative method by Fetzer et al. [10].
For all images, we consider the principal point to lie in their center which is a commonly used assumption. With the exception of the Bougnoux formula, all methods require priors, or initial estimates, for the focal lengths. Following common practice, as established in COLMAP [36], we initialize the priors for focal lengths at $1.2$ times the maximum dimension of the image (i.e., either width or height). This initial setting implies an approximate 50-degree field of view for the cameras.
@@
with $f_{i}^{est}$ and $f_{i}^{gt}$ denoting the estimated and ground truth focal lengths respectively, with $i\in\left\{1,2\right\}$ denoting the first or the second camera. To gauge the accuracy of the poses, we employ the Mean Average Accuracy (mAA) metric, as originally proposed in [17]. The basis for this metric is the pose error $(p_{err})$ which is calculated as the maximum of the rotation and translation error in degrees.
@@ Results:
Hartley [14] PyTorch [30] 659.61 85.38
@@
Fetzer [10] PyTorch [30] 197.89 0.00
@@
Bougnoux [2] NumPy [13] 0.27
@@
Table 2: Mean computation times and number of iterations for the compared methods on the Phototourism dataset [17]. Measurements were performed on Intel i7-11800H. The times shown do not include the estimation of the fundamental matrices.
@@
Hartley [14] 21.51° 12.38 24.26 0.523 0.46 1.12
Fetzer [10] 70.40° 2.31 5.20 0.999 6.86 11.64
Sturm [41] 17.98° 20.64 33.30 0.524 10.78 18.46
Minimal [40] 36.18° 16.42 25.74 0.239 15.04 25.89
@@
Table 3: Mean average accuracy scores for poses (mAAp) and estimated focal lengths (mAAf) on 12 scenes from the ETH3D High Resolution Multiview dataset [37]. The focal lengths of both cameras are assumed to be equal. The last row includes reference results for pose accuracy when ground truth intrinsics are used.
We also evaluate our method for the case when both focal lengths can be assumed to be equal. In this case, we use the modified version of our iterative algorithm from Section 3.1. To perform the evaluation, we use the ETH3D Multiview dataset [37]. We use all image pairs satisfying the co-visibility criterion for each of the 12 scenes from the test set, yielding 3830 total pairs. For comparison, we modified the method by Fetzer et al. to optimize for a single focal length. We use Hartley and Silpa-Anan’s method as usual, but on output we average the two focal lengths produced. Instead of the Bougnoux formula, we use the formula proposed by Sturm [41] which is specific to the case of equal focal lengths. We also evaluate the minimal 6-point algorithm [40] implemented within LO-RANSAC [8, 24]. Further details and additional experiments are available in SM.
The results are shown in Table 3 and Fig. 4(c). Our method provides superior results to other methods based on the decomposition of the fundamental matrix, and performs comparably to the approach based on the minimal solver [40] in terms of estimated focal lengths and shows minor improvements in terms of pose accuracy. Our method thus provides a good alternative to the 6-point minimal solver in the case when the focal lengths of the two cameras are known to be equal.

### v3

@@ 3.2 Real Focal Length Checking within RANSAC
The method proposed in 3.1 is usually applied on a fundamental matrix $\mathbf{F}$ that is obtained using a robust RANSAC-style estimation [ransac, magsac++, lo-ransac, degensac, vsac, prosac]. Compared to the Bougnoux formula [bougnoux_original], our method does not suffer from imaginary focal length estimates in the presence of noise or an error in the principal point location. Thus, our method does not need to check whether the fundamental matrix returned by the RANSAC is decomposible to real focal lengths or not. Nevertheless, as we observed in our experiments, fundamental matrices that result in imaginary focal lengths usually do not provide good estimates and are therefore not selected inside RANSAC as best models.
@@
To reject models, we use a modified version of the Bougnoux formula [bougnoux_original] presented in [oleh_bp], which has the form of a ratio of two polynomials in the elements of the fundamental matrix. The formula returns squared focal lengths. When the sign of one of the squared focal lengths is negative, we reject the model. This approach does not require the use of SVD or other expensive matrix manipulations. This type of checking can be implemented into different variants of RANSACs and, as shown later in our experiments, it leads to reduced computational times while maintaining or even improving accuracy of pose estimates. Implementation details can be found in the Supplementary Material (SM).
@@ 4 Synthetic Experiments
(c) Phototourism [phototourism_benchmark]
Figure 1: Plots showing portion of samples for which our algorithm would converge given a threshold for the relative change of errors in successive iterations $\frac{|e^{n}-e^{n-1}|}{e^{n}}<\epsilon$. For synthetic experiments (a) and (b) we generated 1000 samples with added noise ($\sigma_{n}=1$, $\sigma_{p}=10$). We set the priors as $f_{1}^{p}=660,f_{2}^{p}=440$. For (c) we used the Phototourism benchmark dataset [phototourism_benchmark]. See Section 4 for details.
@@
In all our synthetic and real experiments, we set the weights in (2) to $w_{i}^{f}=5\cdot 10^{-4}$ and $w_{i}^{\mathbf{c}}=1.0$. These weights were obtained by testing different combinations on the Brandenburg Gate validation scene from the Phototurism dataset [phototourism_benchmark] and resulted in the best performance on this scene. However, as we show also in synthetic experiments, for a wide variety of combinations of weights, we consistently obtain very good estimates and outperform the state-of-the-art methods. We set the weights for Hartley and Silpa-Anan’s method [hartley_iterative] in the same manner to 1.0 for $w_{F}$, $10^{-4}$ for $w_{\mathbf{c}}$ and $10^{-6}$ for $w_{f}$.
@@
To perform the experiments, we set up a scene with two cameras. We assume $640~\times~480$ image size and principal points in image centers. The first camera has focal length $f_{1}=600$ and the second $f_{2}=400$. The first camera has center at the origin with the z-axis of the world coordinate system being its principal axis. We perform experiments with the second camera in different configurations denoted as $\mathcal{C}(\theta,y)$. For $\mathcal{C}(0\degree,y)$ the second camera center is located at coordinates $(1200,y,600)$. The camera is rotated by $60\degree$ around its $y$-axis. With $\theta\neq 0\degree$ we additionally rotate the second camera around the $x$-axis by $\theta$. When $y=0$ and $\theta=0\degree$ the principal axes of the two cameras intersect and are thus coplanar, which is a degenerate configuration for the Bougnoux formula [bougnoux_original].
For a given camera setup, we uniformly generate 100 random points visible by both cameras. After projecting the points into both images, we add Gaussian noise with a standard deviation $\sigma_{n}$ to the pixel coordinates. We also emulate errors in the assumed locations of the principal points. The principal points are shifted from the image centers by a distance sampled from the normal distribution with a standard deviation of $\sigma_{p}$. We estimate the fundamental matrix using MAGSAC++ [magsac++] as implemented in OpenCV [opencv].
@@
In the first experiment, we evaluate convergence of the proposed iterative method. Different termination criteria can be considered. We terminate the algorithm when the relative change in error in two successive iterations $\frac{|e^{n}-e^{n-1}|}{e^{n}}$ is less than a predetermined threshold. Fig. 1(a) shows that for non-degenerate configuration our method converges quickly even with strict thresholds. Fig. 1(b) shows the convergence rates of our method when the principal axes are coplanar. The degenerate configuration requires an increased number of iterations to converge with the same thresholds. However, even in a degenerate configuration, the method converges in a few iterations for most of the samples, even with strict thresholds. We also evaluate convergence on a large scale real-world dataset [phototourism_benchmark] (see section 4 for more details). As shown in Fig. 1(c), our method successfully converges for most of the samples even with a strict threshold. Using these findings, we established the maximum number of iterations for our method at 50, as further iterations demonstrated diminishing returns in terms of convergence rate.
@@
In the next experiment, we compare our method with state-of-the-art methods in terms of accuracy of the estimated focal lengths. The methods considered are the baseline Bougnoux formula [bougnoux_original] and the iterative approaches of Hartley and Silpa-Anan [hartley_iterative] and Fetzer et al[fetzer]. In the experiments, we use the same priors for Hartley and Silpa-Anan’s method as for ours. We set the initial values for LM optimization for the method by Fetzer et al. to the same values as our priors.
OursHartley and Silpa-Anan[hartley_iterative]Fetzer et al. [fetzer]Bougnoux[bougnoux_original]GT $f_{1}$Prior $f_{1}$
@@ 5 Real-world Experiments
Figure 3: The plot shows mAAp achieved on the Phototourism datset [phototourism_benchmark] for different RANSAC implementations [opencv, vsac, poselib] with and without real focal length checking (RFC) under varying number of iterations. To calculate the poses we use the ground truth focal lengths. Further details are provided in the SM.
OursHartley and Silpa-Anan[hartley_iterative]Fetzer et al. [fetzer]Bougnoux[bougnoux_original]Sturm [sturm_single]Minimal 6-point [stewenius_single_6pt]PriorRFCOriginal
(a) Phototourism [phototourism_benchmark]
(b) Aachen Day-Night v1.1 [aachen_dn_extended]
(c) ETH3D Multiview [eth3d_multiview]
@@
Phototourism [phototourism_benchmark] Aachen Day-Night v1.1 [aachen_dn_extended]
@@
Table 1: Median errors for poses ($p_{err}$) and focal lengths ($f_{i}^{err}$) and mean average accuracy scores for poses (mAAp) and estimated focal lengths (mAAf) on 12 scenes from the Phototourism dataset [phototourism_benchmark] and the Aachen Day-Night v1.1 dataset [aachen_dn_extended]. RFC denotes real focal length checking as described in subsection 3.2.
@@
We assess the real-world performance of our method using two extensive datasets. The first dataset, referred to as the Phototourism dataset, was originally introduced in [phototourism_benchmark] to serve as a robust benchmark to evaluate structure-from-motion pipelines. The dataset contains 25 scenes of landmarks. Large sets of crowd-sourced images are available for each scene. For 13444Originally, reconstructions were available for 16 of the scenes, but 3 of the scenes were later dropped from the dataset due to data inconsistencies. of the scenes a COLMAP [colmap] reconstruction is provided to serve as ground truth for camera intrinsics and extrinsics. To evaluate the estimated focal lengths, we randomly sample 1000 pairs of images for each scene. We only consider pairs with sufficient co-visibility as defined in [phototourism_benchmark]. We used the Brandenburg Gate scene to serve as a validation set to determine the optimal setting of the weights $w_{i}^{f}$ and $w_{i}^{\mathbf{c}}$ in (2).
The second dataset, known as the Aachen Day-Night v1.1 Dataset, is an extension of the previously established Aachen Day-Night dataset [aachen_dn, aachen_original]. This dataset incorporates a reference COLMAP [colmap] model that has been reconstructed from a comprehensive collection of 6,697 images captured by a variety of cameras and smartphones. For evaluation, we randomly sample 1000 pairs of images that adhere to the same co-visibility criteria as those defined for the Phototourism dataset.
@@
We use LoFTR [loftr] to obtain point correspondences. We perform inference on images resized so that the largest dimension is equal to 1024 px. To estimate the fundamental matrix, we use MAGSAC++ [magsac++] with the epipolar threshold set to 3 px.
We also perform experiments with and without utilizing real focal length checking (RFC) as described in subsection 3.2. Fig 3 shows that on the Phototourism dataset [phototourism_benchmark] the fundamental matrices obtained with RFC generally lead to more accurate pose estimates while also significantly reducing computational times across various RANSAC implementations. Implementation details and further experiments are provided in SM.
@@
We compare our method with three state-of-the-art approaches. We evaluate the Bougnoux formula [bougnoux_original], Hartley and Silpa-Anan’s iterative method [hartley_iterative] and the iterative method by Fetzer et al[fetzer].
For all images, we consider the principal point to lie in their center which is a commonly used assumption. With the exception of the Bougnoux formula, all methods require priors, or initial estimates, for the focal lengths. Following common practice, as established in COLMAP [colmap], we initialize the priors for focal lengths at $1.2$ times the maximum dimension of the image (i.e., either width or height). This initial setting implies an approximate 50-degree field of view for the cameras.
@@
with $f_{i}^{est}$ and $f_{i}^{gt}$ denoting the estimated and ground truth focal lengths respectively, with $i\in\left\{1,2\right\}$ denoting the first or the second camera. To gauge the accuracy of the poses, we employ the Mean Average Accuracy (mAA) metric, as originally proposed in [phototourism_benchmark]. The basis for this metric is the pose error $(p_{err})$ which is calculated as the maximum of the rotation and translation error in degrees.
@@ Results:
Hartley [hartley_iterative] PyTorch [pytorch] 659.61 85.38
@@
Fetzer [fetzer] PyTorch [pytorch] 197.89 0.00
@@
Bougnoux [bougnoux_original] NumPy [numpy] 0.27
@@
Table 2: Mean computation times and number of iterations for the compared methods on the Phototourism dataset [phototourism_benchmark]. Measurements were performed on Intel i7-11800H. The times shown do not include the estimation of the fundamental matrices.
@@
Hartley [hartley_iterative] 21.51° 12.38 24.26 0.523 0.46 1.12
Fetzer [fetzer] 70.40° 2.31 5.20 0.999 6.86 11.64
Sturm [sturm_single] 17.98° 20.64 33.30 0.524 10.78 18.46
Minimal [stewenius_single_6pt] 36.18° 16.42 25.74 0.239 15.04 25.89
@@
Table 3: Mean average accuracy scores for poses (mAAp) and estimated focal lengths (mAAf) on 12 scenes from the ETH3D High Resolution Multiview dataset [eth3d_multiview]. The focal lengths of both cameras are assumed to be equal. The last row includes reference results for pose accuracy when ground truth intrinsics are used.
We also evaluate our method for the case when both focal lengths can be assumed to be equal. In this case, we use the modified version of our iterative algorithm from Section 3.1. To perform the evaluation, we use the ETH3D Multiview dataset [eth3d_multiview]. We use all image pairs satisfying the co-visibility criterion for each of the 12 scenes from the test set, yielding 3830 total pairs. For comparison, we modified the method by Fetzer et al. to optimize for a single focal length. We use Hartley and Silpa-Anan’s method as usual, but on output we average the two focal lengths produced. Instead of the Bougnoux formula, we use the formula proposed by Sturm [sturm_single] which is specific to the case of equal focal lengths. We also evaluate the minimal 6-point algorithm [stewenius_single_6pt] implemented within LO-RANSAC [lo-ransac, poselib]. Further details and additional experiments are available in SM.
The results are shown in Table 3 and Fig. 4(c). Our method provides superior results to other methods based on the decomposition of the fundamental matrix, and performs comparably to the approach based on the minimal solver [stewenius_single_6pt] in terms of estimated focal lengths and shows minor improvements in terms of pose accuracy. Our method thus provides a good alternative to the 6-point minimal solver in the case when the focal lengths of the two cameras are known to be equal.